Two numbers are in the ratio 1:3. If i add 2 to each of the numbers, the ratio becomes 1:2. Find the number. PLease write it in the form of algebraic equation.
Thank You!

Answer by mananth(16949)

Let one number be x
let other number be y
x/y = 1/3
cross multiply 3 x = 1 y
3x -y =0 ...........1
( X +2)/(y+2)= ½
2(x+2)= 1(y+2)2x+4=y+2
2x-y=-2 ............2
Multiply (1) by -1 we get
-3x-y=0 ............3
Add (3)&(2)
-x=-2
x=2
2 & 6